# A systematically-selected sample of luminous, long-duration, ambiguous nuclear transients

### v1

@@ 2.2.2 Spectroscopic follow-up
After visual inspection and spectroscopic confirmation, we retain seven ANTs. We supplement this by searching for events within ZTF that fail our selection but have similar spectroscopic features that do not resemble TDEs or AGN.
@@ 2.3 Supplementary spectroscopic sample
Our method of selecting ANTs with smooth power-law declines excludes events with re-brightening episodes. To supplement this sample, we search for transients that have ANT-like light curves without the smoothness criteria. To do so we search the TNS for events that were observed spectroscopically and classified as AGN or type II superluminous supernovae (SLSNe-II): ANT is not a well-defined spectroscopic class, and BFF is not a classification category in TNS. We find three ZTF-detected events that were not identified by our pipeline but with ANT-like light-curves and spectra: ZTF19aatubsj/AT2019fdr, ZTF20aanxcpf/AT2021loi, and ZTF20acvfraq/AT2020adpi.
@@ 3 A sample of ambiguous nuclear transients
Based on a selection pipeline we identify 59 long-duration, smooth, high-amplitude transient events. Of these, 21 have declines longer than their rises. We classify six as ANTs, one as uncertain, and 14 as extreme-variability AGN. We supplement these seven transients (six ANTs and one Uncertain) with three spectroscopically selected ANTs. In this section we focus on the properties of the 10 selected ANTs, which we denote our ‘main’ sample.
@@ 3.1 Individual objects
We begin here by describing the individual events, before analysing the sample as a whole in Section 3.2. For each object, we check for crossmatches in other surveys: the Asteroid Terrestrial-impact Last Alert System (ATLAS; Tonry et al. 2018; Smith et al. 2020) from which we obtain forced photometry from the ATLAS webserver (Shingles et al., 2021)777https://fallingstar-data.com/forcedphot/; the Panoramic Survey Telescope and Rapid Response System (PanSTARRS; Wainscoat et al. 2016; Chambers et al. 2016; a description of the transient filtering in current operation of the twin Pan-STARRS system is described in Smartt et al. 2024); Gaia Science Alerts (Hodgkin et al., 2021) which are particularly useful for assessing past variability of a given source; the Wide-field Infrared Survey Explorer (WISE) spacecraft as part of the NEOWISE reactivation mission (Mainzer et al., 2011) from the NASA/IPAC infrared science archive (IRSA888https://irsa.ipac.caltech.edu/Missions/wise.html), the Karl G. Jansky Very Large Array Sky Survey (VLASS 2.1; Lacy et al. 2020) via the Canadian Astronomy Data Centre999https://www.cadc-ccda.hia-iha.nrc-cnrc.gc.ca/en/vlass/. We check the archive of the Neil Gehrels Swift Observatory (Swift; Gehrels et al., 2004), and for five events obtained our own target of opportunity observations (PI: Wang). Swift X-ray data were taken in photon-counting mode with the X-ray Telescope (XRT; Burrows et al. 2005) and reduced with the tasks xrtpipeline and xselect. The source and background events were extracted using a circular region of 40 " and an annular ring with inner and outer radii of 60 and 110 ", respectively, both centred at the position of the source. The Ancillary Response Files were created with the task XRTMKARF and the Response Matrix File (RMF), swxpc0to12s6_20130101v014.rmf, was taken from the Calibration Data Base101010https://heasarc.gsfc.nasa.gov/docs/heasarc/caldb/swift. Targets were simultaneously observed with the Ultraviolet-Optical Telescope (UVOT; Roming et al. 2005) in the (central wavelength, 1928 Å), (2246 Å), (2600 Å), (3465 Å), (4392 Å), and (5468 Å) filters. The task UVOTIMSUM to sum all the exposures when more than one snapshot was included in each individual filter data and the task UVOTSOURCE to extract magnitudes from aperture photometry. A circular region of 5 " centred at the target position was chosen for the source event and another region of 40 " located at a nearby position was used to estimate the background emission.
@@
We search the for spectra in the following public archives: the European Southern Observatory (ESO) Science Portal 111111https://archive.eso.org/scienceportal/home, the Keck Observatory Archive 121212https://www2.keck.hawaii.edu/koa/public/koa.php, the Gemini Observatory Archive 131313https://archive.gemini.edu/, the Gran Telescopio CANARIAS Public Archive 141414https://gtc.sdc.cab.inta-csic.es/gtc/index.jsp.
Light curves for the main sample of 10 ANTs are shown in Fig. 3 and spectra are displayed in Fig. 4. Close-ups of the He ii - [O iii]-H region ( Å) and the H region ( Å) is shown in Fig. (5).
@@ 3.1.1 Photometrically selected ANTs
ZTF20abrbeie/AT2021lwx/ATLAS20bkdj/PS22iin, . This event was analysed in Wiseman
@@
et al. (2023). The light curve is smooth and reaches , rising (in the rest frame) for over d and declining smoothly for over d. It remains anomalous amongst the sample presented in this paper in having no detected host galaxy. The spectra displays broad and narrow Balmer emission lines, semi-forbidden carbon C ii] and C iii], and Mg ii. There are no forbidden oxygen lines. A MIR flare is present with a yr rise and subsequent decline. X-rays are detected from the source.
ZTF19aamrjar, . First reported by this work. This object was observed with the Las Cumbres Observatory (LCO Brown et al., 2013) Floyds spectrograph mounted on the 2 m Faulkes Telescope North in Haleakala, USA. The spectrum was reduced using the floyds_pipeline (Valenti et al., 2014)151515https://github.com/LCOGT/floyds_pipeline. It was also observed with the Optical System for Imaging and low-Intermediate-Resolution Integrated Spectroscopy (OSIRIS; Cepa et al. 2000) on the 10.8 m Gran Telescopio Canarias (GTC) at Observatorio del Roque de los Muchachos, La Palma, Spain. The spectrum was acquired in long-slit mode. The data were reduced via the pypeit package (Prochaska et al., 2020b, a). The spectra show strong Balmer and Mg ii emission at . The Balmer lines (H – H) are very broad with full-width at half maximum (FWHM) of km s-1. The line profiles are asymmetric with red wings. [O ii] and [O iii] are strong and narrow. There is Fe ii emission redwards of Å. The spectrum is similar to highly luminous quasars with peculiar line profiles that have been identified as candidate recoiling or binary quasars (Eracleous et al., 2012; Shapovalova et al., 2016).
ZTF19aamrjar has lasted just short of 1000 d in the rest frame since the first detection by ZTF, with a rise and decline timescale greater than that of AT2021lwx. There is a MIR flare evolving slower than the optical, which has lasted 2 yr and not yet reached its peak.
@@
ZTF20abodaps/AT2020afep/ATLAS20vrw, . First reported in this work. No previous variability in PS1 or PTF. The contrast between rise and decline timescales is extreme: the light curve rises to -24 mag in d and declines constantly in 600 d. Observed with LCO Floyds on FTN, and via the European Southern Observatory (ESO) as part of the extended Public ESO Spectroscopic Survey of Transient Objects (ePESSTO+; Smartt
et al. 2015) using the ESO Faint Object Spectrograph and Camera (EFOSC2; Buzzoni
et al. 1984) on the New Technology Telescope (NTT) at ESO La Silla observatory, Chile. The EFOSC2 spectrum was reduced using the PESSTO pipeline (Smartt
@@
et al., 2015) v.3.0.1. The spectrum shows somewhat broadened Balmer lines (H–H) at , along with Mg ii. There is no detection of [O ii] or [O iii], similar to AT2021lwx. Similarly to ZTF19aamrjar, ZTF20abodaps/AT2020afep has a MIR flare evolving slower than the optical, which appears to have reached peak brightness 2 yr after the optical peak. X-ray observations were obtained on MJD 60263.67 ( d) with an exposure of 2367 s. X-rays are marginally detected.We find an upper limit of erg cm-2 s-1 when assuming a power law with index and a foreground absorption of cm-2. This limit corresponds to a luminosity of erg s-1.
@@
ZTF18aczpgwm/AT2019kn/ATLAS19bdfo/Gaia19abv, . First reported by Gaia Science Alerts. The decline rate is very similar to AT2021lwx. A slight rebrightening occurred at +450 d. We obtain two publicly available spectra161616Program IDs C252,C253; PI: Graham from the Keck Low Resolution Imaging Spectrometer (LRIS; Oke et al. 1995) at the W. M. Keck Observatory, Hawaii, USA. The spectra were reduced using the LPIPE pipeline (Perley, 2019) with default settings. The spectrum displays strong emission lines of the Balmer series and Mg ii at , as well as broad forbidden lines of [O iii] and narrow lines of [O ii], O i, and [S ii]. The low-ionization lines (broad and narrow) show asymmetric profiles with a strong blue wing, indicating outflowing material. The presence of strong Fe ii emission indicates a hot, ionizing UV continuum. In addition, high-ionization ‘coronal’ iron lines are present from [Fe vi] and [Fe vii]. The spectrum suggests an underlying AGN similar to a NLSy1. The MIR flare rises for yr, followed by a plateau lasting yr. X-ray observations were obtained on MJDs 60262.82 and 60264.41 ( d, d) with exposures of 1183 s and 2580 s respectively. We find an upper limit of erg cm-2 s-1 when assuming a power law with and an absorption of cm-2. This limit corresponds to a luminosity of erg s-1.
ZTF21abxowzx/AT2021yzu/ATLAS21bjhp, . First reported by ZTF, classified as an AGN (Chu
et al., 2021b). A bona-fide flare, this event occurred in a galaxy with no previous variability and with MIR colours not consistent with AGN activity. We retrieve a Keck/LRIS spectrum from the Weizmann Interactive Supernova Data Repository (WISeREP; Yaron &
@@
Gal-Yam 2012)171717 https://www.wiserep.org/. The spectrum is the bluest of the sample and somewhat resembles a type I AGN (Vanden Berk
et al., 2001). It has broadened Balmer and Mg ii, narrow [O ii] and [O iii], and hints of Fe ii. There is a MIR flare at much lower significance than all the other events studied here.
@@
ZTF20aaqtncr/AT2021fez/ATLAS20pzv/Gaia21bgs, . First reported by Gaia Science Alerts. The light curve has a smooth rise and decline of similar duration. A spectrum was obtained with LCO Floyds FTN. The spectrum shows strong narrow Balmer emission lines, strong [N ii], and very strong [O iii]. The [O iii]/H ratio alone is enough to place the ionizing source as an AGN, which is confirmed by placing the line fluxes in the [O iii]/H vs [N ii/H] Baldwin-Phillips-Terlevich (BPT) diagram (Baldwin
ZTF19aailpwl/AT2019brs/Gaia19axp, . First reported by Gaia Science Alerts. Presented in the ANT sample of Frederick et al. (2021). The light curve is similar to AT2021lwx, although less luminous () and with a faster ( d) rise followed by a plateau and a similar gradient decline. The event occurred in a known NLSy1. The spectrum shows strong narrow Balmer emission along with He ii and N iii characteristic of Bowen fluorescence flares.

### v2

@@ 2.2.2 Spectroscopic history and follow-up
After visual inspection and spectroscopic filtering, we retain seven ANTs which we term our ‘photometrically selected’ sample: ZTF18aczpgwm=AT2019kn, ZTF19aailpwl=AT2019brs, ZTF19aamrjar, ZTF20abodaps=AT2020afep, ZTF20abrbeie=AT2021lwx, ZTF21abxowzx=AT2021yzu, ZTF22aadesap=AT2022fpx. We supplement these seven by searching for events within ZTF that fail our photometric selection but have similar spectroscopic features that do not resemble TDEs or AGN.
@@ 2.3 Supplementary spectroscopic sample
Our method of selecting ANTs with smooth power-law declines excludes genuine transient events with re-brightening episodes. To supplement this sample, we search for transients that have ANT-like light curves without the smoothness criteria. To do so we search the TNS for events that were observed spectroscopically and classified as AGN or type II superluminous supernovae (SLSNe-II) since ANT is not a well-defined spectroscopic class, and BFF is not a classification category in TNS. We inspect the light curves and spectra for ANT-like signatures: a nuclear location, a long-duration ( d) light curve, and a spectrum with a blue continuum and strong Balmer emission lines. We find three ZTF-detected events that were not identified by our pipeline but that fit these criteria. Two of these have been published in the literature as ANTs (ZTF19aatubsj/AT2019fdr and ZTF20aanxcpf/AT2021loi), and one which has not (ZTF20acvfraq/AT2020adpi). We add these to our analysis sample.
@@ 3 A sample of ambiguous nuclear transients
Based on a photometric selection pipeline we identify 58 long-duration, smooth, high-amplitude events. We classify 6 as unclear and 16 as transients, of which five are spectroscopically confirmed supernovae and one is a TDE. Of the remaining nine likely ANTs, seven have spectra that do not resemble supernovae or TDEs. We supplement these seven ANTs with three spectroscopically selected ANTs that have less smoothly evolving light curves. In this section we focus on the properties of these 11 selected ANTs, which we denote our ‘analysis’ sample.
@@ 3.1 Individual objects
We begin here by describing the individual events, before analysing the sample as a whole in Section 3.3. For each object, we check for crossmatches in other surveys. For each survey we describe the data source and reduction methods below. In all cases, we correct for Galactic reddening using the dust maps of Schlafly & Finkbeiner (2011) assuming a Fitzpatrick (1999) extinction law.
3.1.1 ATLAS
We cross-match the ZTF transient with the database of the Asteroid Terrestrial-impact Last Alert System (ATLAS; Tonry et al. 2018; Smith et al. 2020) from which we obtain forced difference imaging photometry from the ATLAS webserver (Shingles et al., 2021)888https://fallingstar-data.com/forcedphot/. Given the slow nature of the ANT light curves, to maximise signal-to-noise we combine ATLAS data into 20-day bins in each of the and filters.
3.1.2 Pan-STARRS
3.1.3 Gaia
We obtain long-term, white-light Gaia -band photometry from via Gaia Science Alerts (Hodgkin et al., 2021). These data are particularly useful for assessing past variability of a given source and are not used for parameter estimation, therefore we do not perform a host galaxy subtraction.
3.1.4 WISE
We obtain mid-infrared photometry from the Wide-field Infrared Survey Explorer (WISE) spacecraft as part of the NEOWISE reactivation mission (Mainzer et al., 2011) from the NASA/IPAC infrared science archive (IRSA999https://irsa.ipac.caltech.edu/Missions/wise.html). We obtain catalogued photometry in the () and () bands for the full duration of the NEOWISE reactivation all sky survey (i.e. since 2013, five years before the start of ZTF). For each transient, we average the flux in each band over the period before the first optical detection and subtract this from post-optical-flare photometry.
3.1.5 VLASS
For each transient we search the Karl G. Jansky Very Large Array Sky Survey (VLASS 2.1; Lacy et al. 2020) via the Canadian Astronomy Data Centre101010https://www.cadc-ccda.hia-iha.nrc-cnrc.gc.ca/en/vlass/.
3.1.6 Swift
@@
We search the archive of the Neil Gehrels Swift Observatory (Swift; Gehrels et al., 2004), and for five events obtained our own target of opportunity observations (PI: Wang). Swift X-ray data were taken in photon-counting mode with the X-ray Telescope (XRT; Burrows et al. 2005) and reduced with the tasks xrtpipeline and xselect. The source and background events were extracted using a circular region of 40 " and an annular ring with inner and outer radii of 60 and 110 ", respectively, both centred at the position of the source. The Ancillary Response Files were created with the task XRTMKARF and the Response Matrix File (RMF), swxpc0to12s6_20130101v014.rmf, was taken from the Calibration Data Base111111https://heasarc.gsfc.nasa.gov/docs/heasarc/caldb/swift. Targets were simultaneously observed with the Ultraviolet-Optical Telescope (UVOT; Roming et al. 2005) in the (central wavelength, 1928 Å), (2246 Å), (2600 Å), (3465 Å), (4392 Å), and (5468 Å) filters. The task UVOTIMSUM to sum all the exposures when more than one snapshot was included in each individual filter data and the task UVOTSOURCE to extract magnitudes from aperture photometry. A circular region of 5 " centred at the target position was chosen for the source event, except for observations later than August 2023 after Swift experienced a jitter, and thus we use 8 " – this affects ZTF18aczpgwm, ZTF19aamrjar, ZTF20abodaps, ZTF21abxowzx. Another region of 40 " located at a nearby position was used to estimate the background emission. We have not performed a host subtraction on the UVOT data. In cases where the UVOT data is used to constrain the black body SED, we verified that fits are consistent with and without the UVOT data.
@@
We search the for spectra in the following public archives: the European Southern Observatory (ESO) Science Portal 121212https://archive.eso.org/scienceportal/home, the Keck Observatory Archive 131313https://www2.keck.hawaii.edu/koa/public/koa.php, the Gemini Observatory Archive 141414https://archive.gemini.edu/, the Gran Telescopio CANARIAS Public Archive 151515https://gtc.sdc.cab.inta-csic.es/gtc/index.jsp.
Light curves for the analysis sample of 8 photometrically selected ANTs are shown in Fig. 3 and the 3 spectroscopically selected are shown in Fig. 4. To convert to absolute magnitude for the figures we include a correction to preserve flux density, but have not -corrected to the rest frame and bands. Spectra are displayed in Fig. 5. Close-ups of the He ii - [O iii]-H region ( Å) and the H region ( Å) is shown in Fig. (6).
@@ 3.1.7 Photometrically selected ANTs
ZTF20abrbeie/AT2021lwx/ATLAS20bkdj/PS22iin, . First reported by ZTF via the AMPEL broker. This event was analysed in Wiseman
@@
et al. (2023). The light curve is smooth and reaches , rising (in the rest frame) for over d and declining smoothly for over d. It remains anomalous amongst the sample presented in this paper in having no detected host galaxy. The spectra displays broad and narrow Balmer emission lines, semi-forbidden carbon C ii] and C iii], and Mg ii. There are no forbidden oxygen lines. A MIR flare is present with a yr rise and subsequent decline. X-rays are detected from the source, consistent with a power-law and inconsistent with extrapolation of the UV-optical black body.
ZTF19aamrjar/ATLAS19mmu, . First reported by this work. ZTF19aamrjar has lasted just short of 1000 d in the rest frame since the first detection by ZTF, with a rise and decline timescale greater than that of AT2021lwx161616Since the initial sample selection and analysis was conducted, ZTF19aamrjar has shown a secondary rise, indicating either more persistent accretion consistent with the quasar-like spectrum or a rebrightening from the transient. Such rebrightenings further blur the line between ANTs and AGN.. There is a MIR flare evolving slower than the optical, which has lasted 2 yr and not yet reached its peak.
This object was observed with the Las Cumbres Observatory (LCO Brown et al., 2013) Floyds spectrograph mounted on the 2 m Faulkes Telescope North in Haleakala, USA. The spectrum was reduced using the floyds_pipeline (Valenti et al., 2014)171717https://github.com/LCOGT/floyds_pipeline. It was also observed with the Optical System for Imaging and low-Intermediate-Resolution Integrated Spectroscopy (OSIRIS; Cepa et al. 2000) on the 10.8 m Gran Telescopio Canarias (GTC) at Observatorio del Roque de los Muchachos, La Palma, Spain. The spectrum was acquired in long-slit mode. The data were reduced via the pypeit package (Prochaska et al., 2020b, a). The spectra show strong Balmer and Mg ii emission at . The Balmer lines (H – H) are very broad with full-width at half maximum (FWHM) of km s-1. The line profiles are asymmetric with red wings. [O ii] and [O iii] are strong and narrow. There is Fe ii emission redwards of Å. The spectrum is similar to highly luminous quasars with peculiar line profiles that have been identified as candidate recoiling or binary quasars (Eracleous et al., 2012; Shapovalova et al., 2016).
@@
ZTF20abodaps/AT2020afep/ATLAS20vrw, . First reported by ePESSTO+. in this work. No previous variability in PS1 or PTF. The contrast between rise and decline timescales is extreme: the light curve rises to -24 mag in d and declines constantly in 600 d. Observed with LCO Floyds on FTN, and via the European Southern Observatory (ESO) as part of the extended Public ESO Spectroscopic Survey of Transient Objects (ePESSTO+; Smartt et al. 2015) using the ESO Faint Object Spectrograph and Camera (EFOSC2; Buzzoni et al. 1984) on the New Technology Telescope (NTT) at ESO La Silla observatory, Chile. The EFOSC2 spectrum was reduced using the PESSTO pipeline (Smartt et al., 2015) v.3.0.1. The spectrum shows somewhat broadened Balmer lines (H–H) at , along with Mg ii. There is no detection of [O ii] or [O iii], similar to AT2021lwx.
Similarly to ZTF19aamrjar, ZTF20abodaps/AT2020afep has a MIR flare evolving slower than the optical, which appears to have reached peak brightness 2 yr after the optical peak.
@@
X-ray observations were obtained on MJD 60263.67 ( d) with an exposure of 2367 s. X-rays are marginally detected.We find an upper limit of erg cm-2 s-1 when assuming a power law with index and a foreground absorption of cm-2. This limit corresponds to a luminosity of erg s-1.
ZTF18aczpgwm/AT2019kn/ATLAS19bdfo/Gaia19abv, . First reported by Gaia Science Alerts. The decline rate is very similar to AT2021lwx. A slight rebrightening occurred at +450 d.
@@
We obtain two publicly available spectra181818Program IDs C252,C253; PI: Graham from the Keck Low Resolution Imaging Spectrometer (LRIS; Oke et al. 1995) at the W. M. Keck Observatory, Hawaii, USA. The spectra were reduced using the LPIPE pipeline (Perley, 2019) with default settings. The spectrum displays strong emission lines of the Balmer series and Mg ii at , as well as broad forbidden lines of [O iii] and narrow lines of [O ii], O i, and [S ii]. The low-ionization lines (broad and narrow) show asymmetric profiles with a strong blue wing, indicating outflowing material. The presence of strong Fe ii emission indicates a hot, ionizing UV continuum. In addition, high-ionization ‘coronal’ iron lines are present from [Fe vi] and [Fe vii]. The spectrum suggests an underlying AGN similar to a NLSy1.
The MIR flare rises for yr, followed by a plateau lasting yr.
X-ray observations were obtained on MJDs 60262.82 and 60264.41 ( d, d) with exposures of 1183 s and 2580 s respectively. We find an upper limit of erg cm-2 s-1 when assuming a power law with and an absorption of cm-2. This limit corresponds to a luminosity of erg s-1.
ZTF21abxowzx/AT2021yzu/ATLAS21bjhp, . First reported by ZTF, classified as an AGN (Chu et al., 2021b). A bona-fide flare, this event occurred in a galaxy with no previous variability and with pre-flare MIR colours (WISE versus ) not consistent with AGN activity.
@@
We retrieve a Keck/LRIS spectrum from the Weizmann Interactive Supernova Data Repository (WISeREP; Yaron & Gal-Yam 2012)191919 https://www.wiserep.org/. The spectrum is the bluest of the sample and somewhat resembles a type I AGN (Vanden Berk et al., 2001). It has broadened Balmer and Mg ii, narrow [O ii] and [O iii], and hints of Fe ii.
There is a MIR flare with a timescale comparable to the optical light curve.
@@
ZTF19aailpwl/AT2019brs/Gaia19axp, . First reported by Gaia Science Alerts. Presented in the ANT sample of Frederick et al. (2021). The light curve is similar to AT2021lwx, although less luminous () and with a faster ( d) rise followed by a plateau and a similar gradient decline.
The event occurred in a known NLSy1. The spectrum (retrieved from WISeREP) shows strong narrow Balmer emission along with He ii and N iii characteristic of Bowen fluorescence flares.
@@
ZTF20abgxlut/AT2020oio/Gaia20dvv/ATLAS20rmk, . First reported by ZTF via the AMPEL broker. Initially classified as an AGN (Terreran et al., 2020), there is no evidence for any variability in pre-existing data.
The transient spectrum (retrieved from WISeREP) is almost identical to ZTF19aailpwl/AT2019brs, except although the He ii, N iii as well as [O iii] are all present, they are all weaker in ZTF20agbxlut.
The pre-flare MIR colours (WISE versus ) are not consistent with AGN activity, and there is a clear MIR flare that lags the optical transient. There are no Swift X-ray observations of this event.
ZTF22aadesap/AT2022fpx/ATLAS22kjn/Gaia22cwy/PS23bdt, . First reported by ATLAS, this is the lowest luminosity event in our sample, showing an initially faster rise and decline followed by a quasi-plateau at later times.
ZTF22aadesap was classified as a TDE by Perez-Fournon et al. (2022) but then identified as an ECLE by Koljonen et al. (2024). While ECLEs have been observed in the late time stages of TDEs (Newsome et al., 2024), the peak spectrum of ZTF22aadesap/AT2022fpx does not resemble the common classes of TDE (e.g. Arcavi et al., 2014).
X-ray observations do not show a signal during the optical peak, but a strong late-time flare is observed at the same time as the optical date plateau. The implications of these observations are discussed in Koljonen
et al. (2024).